Shares in Brothers Grimm, Inc., manufacturers of gingerbread houses, are expected to pay a dividend of $5.00 in one year and to sell for $100 per share at that time. How much should you be willing to pay today per share of Grimm under the following circumstances? Instructions: Enter all responses rounded to two decimal places. a. Consider a safe rate of interest of 5 percent and assume that investing in Grimm carries no risk. Grimm's share value would be $ b. Consider a safe rate of interest of 10 percent and assume that investing in Grimm carries no risk. Grimm's share value would be $ c. Consider a safe rate of interest of 5 percent and a risk premium of 3 percent. Grimm's share value would be $ Now assume that Grimm is not expected to pay a dividend, but the expected price is unchanged
Instructions: Enter all responses rounded to two decimal places. d. Consider a safe rate of interest of 5 percent and assume that investing in Grimm carries no risk Grimm's share value would be$ e. Consider a safe rate of interest of 10 percent and assume that investing in Grimm carries no risk. Grimm's share value would be$ f. Consider a safe rate of interest of 5 percent and a risk premium of 3 percent. Grimm's share value would be$
Answer #1

Solution:-

A):-

The share value today would be

$ 100.00 = (5+100)/(1+5%)

B):-

The share value today would be

$ 95.45 = (5+100)/(1+10%)

C):-

The share value today would be

$ 97.22 = (5+100)/(1+8%)

D):-

The share value today would be

$ 95.24 = 100/(1+5%)

E):-

The share value today would be

$ 90.91 = 100/(1+10%)

F):-

The share value today would be

$ 92.59 = 100/(1+8%)